OpenStudy (anonymous):

x^2-18X=10 solve by completing the square

OpenStudy (anonymous):

use commas or radicals if necessary

OpenStudy (saifoo.khan):

decimals?

OpenStudy (anonymous):

To complete the square take the middle term half it then square it

OpenStudy (anonymous):

Half of -18x is -9 square is 81 x^2 -18x+81=10+81

OpenStudy (anonymous):

± what is this symbol

OpenStudy (saifoo.khan):

+ve & -ve. this means that there will be 2 answers.

OpenStudy (anonymous):

It means + or - both.

OpenStudy (anonymous):

can this answer be in radical form as well?

OpenStudy (anonymous):

it means that for x you will get two values for completing the square first add square of half the coefficient of x to both sides x² - 18x + 9² = 10 + 9² LHS becomes (x - 9)² so (x - 9)² = 91 taking squareroot of both sides x - 9 = root(91) x - 9 = ± root(91) so x = 9 + root(91) and x = 9 - root(91)

OpenStudy (anonymous):

\[x = 9 + \sqrt{91}\] and \[x = 9 - \sqrt{91}\]

OpenStudy (anonymous):

got it panther???

OpenStudy (anonymous):

ok ty